Iodine fluoride gas (IF) decomposes to gaseous iodide (I2) and fluoride (F2) at a high temperature. If Kp at this high temperature is 19.7, determine the equilibrium partial pressure of IF gas if its initial concentration was 0.306 atm? Hint: You should write a balanced chemical equation for the reaction described.

Question 6 options:

0.153 atm

0.0689 atm

0.168 atm

0.0310 atm

0.138 atm
Answer #1

SOLUTION

We have the balanced chemical reaction is

2IF   \rightleftharpoons I2 + F2

Initial concentration   0.306 atm 0 0

equillibrium concentration 0.306 - 2x    x x

Kp = [I2 ][F2]/[IF]2 = 19.7

x*x / ( 0.306 - 2x )2 = 19.7

x2 = 19.7 (0.0937 + 4x2 - 1.224x)

   x2 = 1.845 + 78.8 x2 - 24.11x

77.8 x2  - 24.11x +1.845 = 0

Solving the quadratic equation. we will get the value of x

x = 0.138

Then equillibrium partial pressure of IF = 0.306 - 2x atm = 0.306 - 2* 0.138 = 0.030 atm

So, the correct option is 0.0310 atm Answer